### v1

@@ 4.2 Results
The assessment methodology for SAVN-NAV navigation tasks[28], similar to the HOZ study[33], involves random selection of the agent’s initial position and goal item, with five trials per scenario. Table 2 displays results for all targets and a subset with optimal path lengths above five steps. Upon integrating the LOAT framework into the HOZ system, we observed improvements in both SPL and SR metrics, surpassing previously reproduced results. This underscores LOAT’s effectiveness within graph-based policy networks.
Table 2: Comparison in SAVN-NAV navigation tasks in AI2-THOR.
@@
Results in val spilt of Habitat ObjectNav are shown in Table 3. LOAT improves initial SR of PEANUT. We also analyzed the categorical success rates, as shown in Table 4. The SR for navigating to hard-to-locate small objects (e.g. "plant") has greatly improved.
Table 3: Results in Habitat ObjectNav in Val split.
@@
PEANUT (reproduced) [32] 0.30 0.55
PEANUT w/ LOAT 0.27 0.58
Table 4: Success Rates by Category in Habitat ObjectNav
Method Chair Bed Plant Toilet TV Monitor Sofa
PEANUT (reproduced) [32] 0.78 0.54 0.11 0.58 0.41 0.71
PEANUT w/ LOAT 0.68 0.57 0.32 0.60 0.42 0.64
After analyzing the outcomes from three benchmarks, the LOAT framework demonstrates advantages in three key aspects:
@@
    Offers targeted guidance for navigating towards small, hard-to-locate objects. By leveraging object affinities from both training experiences and LLM insights, it is particularly efficient for detecting objects that are difficult to identify without context from relevant easier-to-find objects.
4.3 Ablation Studies
4.3.1 Impact of the LOAT Framework in ALFRED[22]:
The LOAT-P system dramatically increased ALFRED’s performance in previous experiments. To differentiate the impact of the LOAT framework from other improvements, such as advanced segmentation methods, LOAT is rigorously evaluated by integrating it into the Prompter [10] and FILM [15] systems within the ALFRED benchmark, while keeping all other variables constant. The outcomes, detailed in Table 5, unequivocally show enhancements across all measured metrics upon incorporating the LOAT framework. Notably, the improvement in goal-found rates is a testament to LOAT’s ability to significantly enhance the agent’s proficiency in identifying target objects. This underscores the framework’s effectiveness in augmenting navigational and object-identification capabilities with insights from both LLM-derived object affinities and training-time preferences, thereby contributing to the overall performance uplift.
Table 5: Performance with/without LOAT Framework Integration in ALFRED Benchmark. We integrate the LOAT framework into FILM[15] and Prompter[10] systems and witness consistent improvements across all metrics.
@@ 4.3.1 Impact of the LOAT Framework in ALFRED[22]:
4.3.2 Impact of LOAT Submodules:
The LOAT framework’s integration of Generalized Affinities (G.A.) and Experiential Affinities (E.A.) modules plays a pivotal role in enhancing navigation and task execution, as evidenced by the ablation study in Table 6. The G.A. module slightly improves success rates, leveraging general knowledge to guide decision-making, but its impact is more nuanced in complex scenarios, suggesting its limitations without learned experience from training environments. Conversely, the E.A. module manages to improve the outcomes in both short and long trials by applying specific past experiences.
@@ 4.3.2 Impact of LOAT Submodules:
Table 6: Comparative evaluation of HOZ incorporating LOAT framework variants, Experiential Affinities (E.A.) and Generalized Affinities (G.A.) modules, indicating performance changes against the HOZ baseline in SAVN-NAV navigation tasks[28].
@@
w/ Full LOAT 39.56 normal-$\uparrow$ 73.12 normal-$\uparrow$ 39.68 normal-$\uparrow$ 65.26 normal-$\uparrow$
In addition to leveraging commonsense from LLMs and experiential object affinities in a parallel manner, we further explore integrating LLM commonsense by employing it as a constraint. Further details can be found in supplementary materials.
4.4 Attention Visualization
@@
We also investigate the attention transition of models across time steps. As shown in Fig.  4, model with only experiential affinities module tends to focus on fixed areas while model with full integration with LOAT is able to adjust focus areas over time. Initially honed in on specific features, the attention gradually broadens to incorporate additional relevant features as dictated by LLM guidance when initial attempts fail to locate the target, showcasing adaptive attention refinement over time.
Figure 4: Predicted Attention Shifts across Time Steps. A model without Generalized Affinties Module tends to fixate on certain areas, while one incorporating full version of LOAT can adapt its focus over time. Initially targeting specific features, attention gradually expands to include more relevant ones based on LLM guidance, showcasing adaptive refinement capability of LOAT.
4.5 Evaluation on Out-of-Domain Objects
The effectiveness of our model in recognizing out-of-domain objects was assessed using 300 pre-collected semantic maps from AI2-THOR environments. We aimed to predict potential locations for unseen target objects based on their category names, involving "Umbrella", "HairDrier", "Scissors", "Toothbrush", "Comb", "Peach", "CanOpener", "Whisk", "Magazine", "Eyeglasses". The model attempts to identify the nearest objects stored in the map to these targets, applying a distance threshold set at one-fifteenth of the map’s total resolution for this experiment. Our primary focus was on assessing the semantic logic behind these probability distributions to determine if the predicted locations for the target objects were plausible within a domestic context. Fig. 5 shows that models relying solely on the experience transfer module may predict a uniform distribution, resulting in semantically incongruous predictions (e.g. associating a can opener with a bathtub basin, which lacks semantic plausibility). LOAT-integrated models better incorporate commonsense reasoning from large language models, resulting in more logically coherent probability distributions (e.g. placing a can opener near cabinets or coffee tables) reflecting intuitive home placement expectations.
@@ 4.5 Evaluation on Out-of-Domain Objects
5 Conclusions
@@
Future research could extend LOAT’s dual-module strategy to open vocabulary navigation, exploring its potential to dynamically adapt to an even broader range of environments and tasks, further pushing the boundaries of autonomous navigation.
Advancing Object Goal Navigation Through LLM-enhanced Object Affinities Transfer - Supplementary Material
Mengying Lin Yaran Chen* Dongbin Zhao Zhaoran Wang
The supplementary material includes an extended ablation study on the dynamic fusion module (Section 0.A), an exploration of using LLM commonsense knowledge as a constraint instead of a parallel information source (Section 0.B), additional examples of attention visualization (Section 0.C), and detailed descriptions of the semantic map formats used in our experiments (Section 0.D).
Appendix 0.A Ablation on Dynamic Fusion Module
We conducted experiments within ALFRED by replacing the dynamic fusion module with fixed guidance rates of 0, 0.5, and 1, with the outcomes detailed in Table 7. For clarify, a guidance rate of 0 corresponds to scenarios where the generalized affinities module is omitted, whereas a guidance rate of 1 signifies the exclusion of the experiential affinities module.
Table 7: Performance with Different Guidance Rate Settings in ALFRED Benchmark. Model having a dynamic fusion module is denoted by Dynamic, while the rest are models with fixed guidance rates.
Guidance Rate Valid Seen Valid Unseen
SR GFR SR GFR
0 61.22 91.83 62.00 87.21
0.5 62.71 92.55 59.95 87.42
1 60.37 92.07 60.41 87.58
Dynamic 59.15 92.56 63.22 90.13
Analyzing the experimental results shown in Table 7, we observe that the dynamic fusion module distinctly enhances the GFR, especially in unseen environments. This emphasizes the module’s ability to effectively assist in locating navigation goals, highlighting the importance and effectiveness of dynamic adaptation in processing LLM insights and historical affinities.
With fixed guidance rates, we see that while SR might fluctuate, the dynamic fusion approach consistently achieves higher GFR, indicating superior performance in identifying target objects even if it does not always translate into higher SR due to possible interaction execution failures within the AI2-THOR simulator. Particularly in unseen settings, the model with dynamic fusion outperforms the fixed-rate configurations, suggesting its superior capacity to generalize and apply learned and inferred object affinities to novel situations.
Appendix 0.B Comparative Analysis of LLM Integration Strategies for LOAT
Building on the previously introduced method, we further explore integrating LLM commonsense by employing it as a constraint, differing from its parallel use as an information source. This exploration led to the development of two distinct designs: LOAT-AVG, the design we use throughout previous sections, where experiences from both sources are averaged, and the newly introduced LOAT-MUL, which uses a multiplicative attention mechanism. In LOAT-MUL, embeddings for target and environmental objects serve as queries and keys. The values, determined by the semantic proximity as per the prior knowledge of large language models, are assigned such that objects semantically close to the target are emphasized (assigned a value of 1), while others are de-emphasized (assigned a value of 0).
The mathematical formulation is as follows, employing the same notation introduced in the Method section. We derive key queries by transforming embeddings of the target and environmental object categories. The values are the semantic relevance of each object $o_{i}$ in the map to the target $o_{target}$, repesented by a binary value $S(o_{i},o_{target})$ determined by the LLM. The output of the LOAT-MUL attention mechanism, $A_{MUL_{i}}$ for each object $o_{i}$, is then calculated as:
$A_{MUL_{i}}=S(o_{i},o_{target})\cdot\frac{\exp(\mathbf{Q}\cdot\mathbf{K}^{T}/\sqrt{d_{k}})}{\sum_{j=1}^{M}\exp(\mathbf{Q}\cdot\mathbf{K}_{j}^{T}/\sqrt{d_{k}})}$ (8)
This method uses LLM knowledge as a gating mechanism to refine attention by emphasizing objects that are closely connected to the objective and ignoring all other objects that are judged irrelevant. Comparative results between the two LOAT designs in the ALFRED benchmark are detailed in Table 8. LLM-AVG-P exhibits higher goal found rates (GFR) in both seen and unseen scenarios compared to LOAT-MUL-P, despite achieving lower success rates (SR). This discrepancy may result from interaction execution failures (e.g. Fail to pick up objects.) within the AI2-THOR simulator environment. While performance between the two methods is comparable, LLM-AVG-P demonstrates a slight advantage in target object localization. This advantage is attributed to potential oversight of objects with subtle but significant associations in terms of LOAT-MUL-P, owing to its gating mechanism. Conversely, LLM-AVG-P retains a more nuanced understanding of object relationships by averaging direct object affinities and LLM guidance, ensuring more inclusive attention allocation.
Table 8: Performance with Different LOAT Framework Designs in ALFRED Benchmark. LOAT-AVG-P is the LOAT-P referred in previous context. LOAT-MUL-P uses LOAT-MUL to replace the original LOAT-AVG design in the LOAT-P system.
Method Valid Seen Valid Unseen
SR GFR SR GFR
LOAT-MUL-P 59.51 91.83 63.93 90.01
LOAT-AVG-P 59.15 92.56 63.22 90.13
Appendix 0.C More Attention Visualization Examples
We also utilize the same approach as described in Attention Visualization section to display the attention for models with LOAT-MUL. As mentioned previously, the attention distribution in the FILM model is notably uniform, lacking focused areas that signify object preference. However, the LOAT framework significantly sharpens models’ focus on objects with semantic relevance, reducing attention to extraneous details. This section delves deeper into the comparison between LOAT-MUL and LOAT-AVG models. As depicted in Fig. 6, LOAT-MUL models employ LLM-derived commonsense knowledge as a stringent filter and achieve a more focused attention. However, it could be argued that there is a tendency for overly selective attention distribution, potentially leading to the oversight of objects with subtle yet valuable associations. For the LOAT-AVG models, which synthesize insights from both commonsense knowledge and experience transfer through a weighted average, exhibit a more balanced attention distribution. This approach ensures a nuanced understanding, retaining objects with marginal relevance learned during training. This balance allows for the preservation of potentially significant objects that might not be immediately obvious, providing a more comprehensive attention strategy.
Figure 6: Attention Visualization Examples. LOAT-MUL models have more selective focus compared with LOAT-AVG models.
Appendix 0.D Semantic Map Format Details
This section outlines the semantic map formats utilized in our navigation task experiments across three benchmarks:
ALFRED[22] Benchmark: Our approach adopts the FILM[15] model setup, leveraging metric maps with a shape of (24, 240, 240). This configuration signifies the inclusion of 24 object categories within a 240x240 resolution map.
SAVN-NAV[28] Navigation Task: Building on the HOZ[33] framework, our model integrates hierarchical object-to-zone graphs encompassing object, zone, and scene layers. Specifically, in room-wise HOZ graphs, zone nodes are derived from clustering egocentric observation features, with edges representing the adjacency likelihood between zones—this probability echoes the co-occurrence rate of objects within zones. Subsequently, room-wise HOZ graphs are amalgamated into comprehensive scene-wise graphs for a holistic environmental depiction.
Habitat ObjectNav[29] Benchmark: Aligning with the PEANUT[32] model, our system employs metric maps of shape (9, 960, 960), which account for 9 object categories within a high-resolution 960x960 map.

### v2

@@ IV-B Results
The experiment results in ALFRED are shown in Table II. LOAT-P achieves SOTA performance across all metrics, with a notable 10% increase in SR. Notably, the SR discrepancy between familiar and unfamiliar environments is markedly reduced, showcasing LOAT’s efficacy in applying generalized object affinities from LLM in novel settings. This underscores the system’s robust adaptability and effectiveness across diverse scenarios.
The assessment methodology for SAVN-NAV navigation tasks[27], similar to the HOZ study[4], involves random selection of the agent’s initial position and goal item, with five trials per scenario. Table III displays results for all targets and a subset with optimal path lengths above five steps. Upon integrating the LOAT framework into the HOZ system, we observed improvements in both SPL and SR metrics, surpassing previously reproduced results. This underscores LOAT’s effectiveness within graph-based policy networks.
TABLE III: Comparison in SAVN-NAV navigation tasks in AI2-THOR.
@@
TABLE IV: Results in Habitat ObjectNav in Val split.
@@
Habitat-Web [41] 0.22 0.55
ProcTHOR [42] 0.32 0.54
PIRLNav[43] 0.28 0.62
PEANUT (reproduced) 0.30 0.55
PEANUT w/ LOAT 0.32 0.63
PEANUT (paper)  [^2] We have confirmed with the author that the validation results in the paper are evaluated on 500 episodes from validation set. See this link https://github.com/ajzhai/PEANUT/issues/3. [30] 0.32 0.64
Results in validation spilt of Habitat ObjectNav are shown in Table IV. LOAT improves the reproduced results of PEANUT.
Comprehensive analysis across three benchmark evaluations reveals that the LOAT framework excels in three key aspects:
@@
    Offers targeted guidance for navigating towards small, hard-to-locate or unseen objects. By leveraging object affinities from both training experiences and LLM insights, it is particularly efficient for detecting objects that are difficult to identify without context from relevant easier-to-find objects.
IV-C Ablation Studies
IV-C1 Impact of the LOAT Framework in ALFRED[26]
The LOAT-P system dramatically increased ALFRED’s performance in previous experiments. To differentiate the impact of the LOAT framework from other improvements, such as advanced segmentation methods, LOAT is rigorously evaluated by integrating it into the Prompter [23] and FILM [32] systems within the ALFRED benchmark, while keeping all other variables constant. The outcomes, detailed in Table V, unequivocally show enhancements across all measured metrics upon incorporating the LOAT framework. Notably, the improvement in goal-found rates is a testament to LOAT’s ability to significantly enhance the agent’s proficiency in identifying target objects. This underscores the framework’s effectiveness in augmenting navigational and object-identification capabilities with insights from both LLM-derived object affinities and training-time preferences, thereby contributing to the overall performance uplift.
Figure 3: Predicted Attention Shifts across Time Steps. A model without the Generalized Affinities Module tends to fixate on certain areas—such as dining tables when searching for a bowl. In contrast, a model incorporating the full version of LOAT can adapt its focus over time, redirecting attention to other meaningful regions like countertops and sinks.
TABLE V: Performance with/without LOAT Framework Integration in ALFRED Benchmark. We integrate the LOAT framework into FILM[32] and Prompter[23] systems and witness consistent improvements across all metrics.
@@ IV-C1 Impact of the LOAT Framework in ALFRED[26]
IV-C2 Impact of LOAT Submodules
The LOAT framework’s integration of Generalized Affinities (G.A.) and Experiential Affinities (E.A.) modules plays a pivotal role in enhancing navigation and task execution, as evidenced by the ablation study in Table VI. The G.A. module slightly improves success rates, leveraging general knowledge to guide decision-making, but its impact is more nuanced in complex scenarios, suggesting its limitations without learned experience from training environments. Conversely, the E.A. module manages to improve the outcomes in both short and long trials by applying specific past experiences. Additionally, an ablation study of the dynamic fusion module (see supplementary) confirms its effectiveness in adaptively incorporating LLM insights and experiential affinities for goal localization.
@@ IV-C2 Impact of LOAT Submodules
TABLE VI: Comparative evaluation of HOZ incorporating LOAT framework variants, Experiential Affinities (E.A.) and Generalized Affinities (G.A.) modules, indicating performance changes against the HOZ baseline in SAVN-NAV navigation tasks[27].
@@
w/ Full LOAT 39.56 $\uparrow$ 73.12 $\uparrow$ 39.68 $\uparrow$ 65.26 $\uparrow$
In addition to leveraging commonsense from LLMs and experiential object affinities in a parallel manner, we further explore integrating LLM commonsense by employing it as a constraint. Further details can be found in the supplementary materials.
IV-D Attention Visualization
@@
We also investigate the attention transition of models across time steps. As shown in Fig.  3, a model without the Generalized Affinities Module tends to fixate on specific areas, such as dining tables when searching for a bowl. In contrast, a model fully integrated with LOAT exhibits the ability to adjust its focus over time. Initially concentrated on particular features, the model’s attention gradually expands to include additional relevant areas, considering countertops and sinks as guided by LLM feedback. This adaptive attention refinement is particularly beneficial when initial attempts to locate the target fail. It underscores the model’s capacity to dynamically adjust and optimize its search strategy over time.
IV-E Evaluation on Out-of-Domain Objects
By utilizing text embeddings instead of one-hot encodings for task representation, we can further examine the model performance over unseen targets in a zero shot manner. The effectiveness of our model in recognizing out-of-domain objects was assessed using 300 pre-collected semantic maps from AI2-THOR environments. We aimed to predict potential locations for unseen target objects based on their category names, involving ”Umbrella”, ”HairDrier”, ”Scissors”, ”Toothbrush”, ”Comb”, ”Peach”, ”CanOpener”, ”Whisk”, ”Magazine”, ”Eyeglasses”. The model attempts to identify the nearest objects stored in the map to these targets, applying a distance threshold set at one-fifteenth of the map’s total resolution for this experiment. Our primary focus was on assessing the semantic logic behind these probability distributions to determine if the predicted locations for the target objects were plausible within a domestic context. Fig. 5 shows that models relying solely on the experience transfer module may predict a uniform distribution, resulting in semantically incongruous predictions (e.g. associating a can opener with a bathtub basin, which lacks semantic plausibility). LOAT-integrated models better incorporate commonsense reasoning from LLMs, resulting in more logically coherent probability distributions (e.g. placing a can opener near cabinets or coffee tables) reflecting intuitive home placement expectations.
@@ IV-E Evaluation on Out-of-Domain Objects
IV-F Navigation in Real Environments
Figure 6: Real-world Navigation Examples. In real world experiments, model with LOAT was able to make reasonable predictions without extra training.
We deployed the LOAT-P navigation system in two scenes adapted from lab spaces, each designed to represent minimal household environments, featuring typical household objects with clear semantic divisions corresponding to different room types. The system was tasked with searching for five object categories—Apple, Book, Bowl, and Cup, Laptop—performing three runs per scene with varying initial positions. Transferred directly from simulation (trained on the ALFRED dataset) to the real world without additional training, the system was deployed on the LoCoBot wx250s. Fig. 6 provides visual examples of the prediction results, demonstrating that the LOAT system is capable of making reasonable predictions in real-world environments. For instance, when searching for a bowl, the predicted goal position was near the microwave.
In all instances, the robot successfully located the target objects, showcasing the LOAT system’s ability to generalize effectively to previously unseen environments.
V Conclusions
@@
While LOAT’s performance can be influenced by the diversity of object categories in the underlying semantic map, future research could explore extending LOAT’s dual-module strategy to open vocabulary navigation. Such an approach holds promise for expanding its adaptability to a broader range of environments and tasks, further advancing autonomous navigation capabilities.